# PAMS: The Perseus Arm Molecular Survey – I. Survey description and first results

### v1

@@ 4.2 Molecular cloud properties
In this Section, we explore some of the basic properties of molecular cloud structures within PAMS, and compare those to an Inner Galaxy reference from the CHIMPS survey (Rigby et al., 2019). The reference sample was restricted to include only those clouds with distances between 2–4 kpc which approximately matches the range in distance of the PAMS sources and thus limits the effect of distance biases. Since CHIMPS covers a longitude range of , the distance limitation results in the Inner Galaxy sample covering a range in Galactocentric radius of kpc, compared to roughly 9–10 kpc for the PAMS Outer Galaxy sample. We also applied a minimum peak SNR criterion to all of our catalogues in order to emulate the ‘reliability’ flagging that was made in the CHIMPS catalogue. In the CHIMPS catalogue, 95% of sources with the highest-reliability flags have a peak SNR greater than 9 and, similarly, 95% of sources with the lowest reliability flag have a peak SNR less than 9, and so we adopt this value for our cut. This conservative cut helps eliminates potentially spurious sources that FellWalker can produce at low SNR, which often appear as separated islands of low-intensity emission (which we refer to as ‘archipelagos’). We adopt the same cut for the CHIMPS Inner Galaxy sample as opposed to using the reliability flags for consistency. By applying these cuts the fwhires, fwlores, fwchimps, and Inner Galaxy catalogues were reduced from 603, 377, 147, and 4999 sources to 265, 54, 76, and 86 sources, respectively.
@@ 4.2.1 Source radii
There are several ways of reporting the size of molecular clouds, which each have drawbacks due to the intrinsic difficulty of representing the complex morphologies and intensity distributions with simple metrics that can be given in a catalogue. These differences may have implications for the way in which key scaling relationships, such as the size–linewidth relationship (Larson, 1981), are compared between different data sets and so we explore them briefly here.
@@
where is the source distance and and are the intensity-weighted rms dispersions in the - and -axes of the image (in the case of PAMS, R.A. and Dec., respectively), deconvolved by the effective beam size. In the case of a perfect Gaussian source detected at high signal-to-noise ratio, would give the standard deviation of the source profile, equivalent to the FWHM / . An alternative is the radius of a circle with the equivalent angular area, , of the source:
@@
to capture the ratio of the two measurements. These two radii differ in approach because depends upon the intensity distribution of the source, while depends only upon the area of the footprint of the source. A cloud with a compact and bright centre surrounded by diffuse emission will, therefore, have a much smaller value of than . is more easily impacted by the sensitivity of the observations, and will recover larger values in deeper observations that detect more diffuse emission. By contrast is more reproducible by observations of different sensitivity and so we generally favour this prescription (and indeed is weakly correlated with peak SNR). Making both measurements of the radius will allow maximum compatibility with other measurements in the literature which use either method, and environmental trends may also reveal themselves in the relationship between these two measurements.
In Fig. 8 we show the relationship between and for our various FellWalker source extractions along with their distributions, and with a comparison to the values reported by Rigby et al. (2019) for the Inner Galaxy from the distance-limited CHIMPS sample. The much larger CHIMPS sample is illustrated as a two-dimensional hexagonal histogram to allow the point density to be seen more easily. In all cases, we find that the relationship between and is non-linear, and is well fitted by a power law. We performed a power law fit to the two radii types for each sample using an orthogonal distance regression444Using scipy.odr. to account for the uncertainties on both variables. The fit results are reported in Table 5. We find that the fwhires and fwlores source extractions produce essentially identical relationships, with power law indices of and , respectively. This is unsurprising because the two extractions are very similar, differing only in the level of fragmentation that they allow (fwhires exist within fwlores structures).
@@
| fwhires | / pc | / pc | 0.270 | 0.009 | 0.692 | 0.013 |
| fwlores | / pc | / pc | 0.275 | 0.018 | 0.707 | 0.027 |
| fwchimps | / pc | / pc | 0.205 | 0.016 | 0.822 | 0.028 |
@@
| fwhires | / pc | / km s-1 | 0.066 | 0.015 | 0.393 | 0.030 |
| fwlores | / pc | / km s-1 | -0.041 | 0.019 | 0.424 | 0.031 |
| fwchimps | / pc | / km s-1 | 0.033 | 0.021 | 0.458 | 0.051 |
@@
| fwhires | / pc | / M⊙ | 3.264 | 0.038 | 2.054 | 0.068 |
| fwlores | / pc | / M⊙ | 2.977 | 0.061 | 1.820 | 0.101 |
| fwchimps | / pc | / M⊙ | 3.514 | 0.061 | 2.136 | 0.119 |
The relationship between the two radius types for the fwchimps and the Inner Galaxy samples are also non-linear and with power law indices of and , respectively, and are consistent with each other. This suggests that there are no significant differences in the mean emission profile for sources at Galactocentric distances of 9–10 kpc compared to sources at 5–7 kpc.
The difference between the fit to the fwchimps data and the fwlores and fwhires is caused by the differences in data quality and to the FellWalker parameter setup, which we have modified in PAMS compared to CHIMPS. We note that some of the non-fitted points in Figs. 8a and b show unusually high or low values of . In cases where is much larger than , these are the archipelago sources that FellWalker identifies at low SNR. The sources with much larger than look like well-recovered sources that are diffuse and have flat emission profiles, and these are over-represented in the data due to a selection bias in FellWalker; the requirement for sources to have a minimum number of pixels above the intensity defined as the noise level prefers diffuse over compact sources. Such sources are also more likely to be the beneficiary of flux boosting effects, in which positive contributions to the emission from the noise may be represented in the data, but negative contributions will not.
Rigby et al. (2019) reported a median value of across the full CHIMPS sample, and the same value for our distance-limited Inner Galaxy sample (Fig. 8d). For fwchimps, the figure is slightly larger at 2.1. By contrast, the fwhires and fwlores median values are larger at 2.7 and 2.3, respectively, which is what is expected for more sensitive data. Overall, the emission profiles of the PAMS sources are similar to those of the Inner Galaxy sample, suggesting that any differences in emission characteristics between 5–7 and 9–10 kpc are mild.
@@ 4.2.2 Scaling relationships
where is the molecular weight per hydrogen molecule, with a value of 2.8 (accounting for a 71% abundance of hydrogen, 27% helium, and 2% metals), is the mass of a hydrogen atom, is the source distance, is the X-factor calculated for 13CO (3–2) derived in Section 4.1, is the integrated intensity of 13CO (3–2) per pixel, and d is the angular area of a pixel. We adopted the method i) value of = cm-2 (km s-1)-1 as our best overall estimate for resolved regions, with a multiplicative factor of 1.5 as its uncertainty. We note that although Rigby et al. (2019) calculated clump masses for the CHIMPS sources using an LTE analysis, we apply the same derivation here for maximum consistency with the PAMS data. We will calculate the masses of PAMS sources using LTE analysis for a comparison with the CHIMPS LTE masses and excitation conditions in a future paper.
@@
We fitted power laws to the size–linewidth and mass–radius relationships using orthogonal distance regression, as in Section 4.2.1, and present the results in Table 5. Fig. 9a shows that the size–linewidth relationships for the fwlores and fwhires extractions follow the same relationship, with the same fitted power law, differing only in a larger normalisation value for the fwhires sources. This is a result of the greater level of fragmentation allowed in the fwhires extraction, where larger linewidths are recovered due to intra-cloud variations that are averaged out in the corresponding fwlores extraction. This is illustrated by the lines connecting the fwhires sources to their fwlores parent. For example, the NGC7538 region catalogues contain 58 and 132 entries in the fwlores and fwhires extractions, respectively, but the bulk of the mass (95% of the 13CO emission) is recovered as a single source in the fwlores catalogue, which is the parent structure of 71 of the sources (i.e. around half) featuring in the fwhires catalogue. The differences between the fwchimps and CHIMPS Inner Galaxy size–linewidth relationships are more different, but are still not significant, given the uncertainties on the fitted parameters.
@@
The mass–radius relationships show more significant differences. Comparing the fwlores and fwhires source extractions, the fwhires sources have a steeper power-law index of 2.05 compared to 1.82 to fwlores, but again the difference is not significantly different when considering the uncertainties. The weak differences in the distributions correspond to a scale dependence in the density distributions of the sources; we note the greater frequency of clump-scale substructures – with sizes of pc and masses of M⊙– in the fwhires distribution. When comparing the fwchimps and Inner Galaxy samples, the power-law index is steeper in the Inner Galaxy compared to the Outer Galaxy, with a value of 2.49 for CHIMPS compared to 2.14 for PAMS, but they are consistent within the uncertainties. This consistency indicates that the mean density profile of the sources are also similar. Surprisingly, the sources from the PAMS region are more massive than their Inner Galaxy counterparts at a given size scale.
@@
In Fig. 9, we also illustrate the relationship of (Kauffmann & Pillai, 2010) that delineates the parameter space that is found, empirically, to be devoid from high-mass star formation in nearby clouds, adapted to the dust opacities used by BGPS (Dunham et al., 2011) and ATLASGAL (Urquhart et al., 2014, 2018). This demonstrates that around half of the sources in PAMS are capable of forming high-mass stars. The proportion appears to be higher, in fact, than is the case for the CHIMPS survey, and we will explore why this might be the case in Section 5.
@@ 4.3 Galactocentric dependence
The combination of the PAMS data with those of CHIMPS allows for an expanded study of properties as a function of Galactocentric distance. The longitude coverage of CHIMPS means that the only clouds in the survey that lie outside the solar circle are at the far side of the Galaxy, with distances in the range 12–17 kpc. Consequently, those clouds are both few in number, and sample only the extreme high-column density (and therefore high-mass) end of the underlying distributions due to Malmquist bias. The PAMS data, therefore, make an important contribution to Galactocentric trends by significantly improving the population statistics at 9–10 kpc, with much improved spatial resolution.
@@
In Fig. 10a we display the mass distributions from CHIMPS (grayscale hexagonal histogram) and PAMS (blue hexagonal histogram). For both PAMS and CHIMPS, we take masses calculated from our factor from Section 4.1. We determined mean radial trends on a subsample of the data, by first excluding all sources with peak SNRs of less than 9, as in Sections 4.2.1 and LABEL:sec:sclaing, and we also applied the mass completeness threshold derived by Rigby et al. (2019, their equation B.1) – shown as the red dashed line – removing all sources at kpc, and all sources with M⊙. The completeness threshold should remove the effects of Malmquist bias. The solid orange line in Fig. 10 shows the moving average value of , with a window size of 0.1 kpc, to all sources in the subsample. It is apparent that the PAMS clumps are consistent with their Inner Galaxy counterparts, and we do not see any systematic trends over the range of probed. This is not affected by the different FellWalker parameter configurations. As far as we can tell with these data, the cloud mass distributions in 13CO (3–2) do not vary with Galactocentric distance.
@@
In Section 4.2.2, we found differences in the power law index of the mass–radius relationship between PAMS and CHIMPS clouds, indicating that the two populations may have different density profiles.We therefore examine the distribution of average volume densities in Fig. 10b, calculated by:
We calculated average radial profiles using the same reduced sample above containing only sources satisfying the minimum mass and SNR criteria which are, again, shown in orange. The distributions of PAMS densities for the fwlores and fwhires extractions show no appreciable difference when compared to the CHIMPS densities. However, the PAMS densities within the fwchimps distribution are around an order of magnitude higher than their CHIMPS counterparts, which is somewhat surprising, and we see the same offset when considering the mean radial trend. The most likely explanation as to why we see this in the fwchimps sources, but not the fwlores and fwhires sources, is that our the factors used for the mass determination are better suited to the latter extractions. We explore this in more detail in 5.

### v3

@@ 4.2 Molecular-cloud properties
In this Section, we explore some of the basic properties of molecular-cloud structures within PAMS, and compare those with an Inner Galaxy reference from the CHIMPS survey (Rigby et al., 2019). The reference sample was restricted to include only those clouds with distances between 2–4 kpc which approximately matches the range in distance of the PAMS sources and thus limits the effect of distance biases. Since CHIMPS covers a longitude range of , the distance limitation results in the Inner Galaxy sample covering a range in Galactocentric radius of kpc, compared with roughly 9–10 kpc for the PAMS Outer Galaxy sample. We also applied a minimum peak SNR criterion to all of our catalogues in order to emulate the ‘reliability’ flagging that was made in the CHIMPS catalogue. In the CHIMPS catalogue, 95 per cent of sources with the highest-reliability flags have a peak SNR greater than 9 and, similarly, 95 per cent of sources with the lowest reliability flag have a peak SNR less than 9, and so we adopt this value for our cut. This conservative cut helps eliminate potentially spurious sources that FellWalker can produce at low SNR, which often appear as separated islands of low-intensity emission (which we refer to as ‘archipelagos’). We adopt the same cut for the CHIMPS Inner Galaxy sample (as opposed to using the reliability flags directly) for consistency. By applying these cuts, the fwhires, fwlores, fwchimps, and Inner Galaxy catalogues were reduced from 948, 483, 231, and 4999 sources to 499, 80, 140, and 865 sources, respectively.
@@ 4.2.1 Source radii
There are several ways of reporting the size of molecular clouds, which each have drawbacks due to the intrinsic difficulty of representing the complex morphologies and intensity distributions with simple metrics that can be given in a catalogue. These differences may have implications for the way in which key scaling relationships, such as the size–linewidth relationship (Larson, 1981), are compared between different data sets and so we explore two of them briefly here.
@@
where is the source distance and and are the intensity-weighted rms dispersions in the - and -axes of the image (in the case of PAMS, R.A. and Dec., respectively), deconvolved by the effective beam size. In the case of a perfect Gaussian source detected at a high signal-to-noise ratio, would give the standard deviation of the source profile, equivalent to the FWHM / . An alternative is the radius of a circle with the equivalent angular area, , of the source. The equivalent radius is:
@@
to capture the ratio of the two measurements. These two radii differ in approach because depends upon the intensity distribution of the source, while depends only upon the area of the footprint of the source. A cloud with a compact and bright centre surrounded by diffuse emission will, therefore, have a much smaller value of than . is more easily impacted by the sensitivity of the observations, and will recover larger values in deeper observations that detect more diffuse emission. By contrast varies less across observations of different sensitivity and so we generally favour this prescription (and indeed is weakly correlated with peak SNR). Making both measurements of the radius will allow maximum compatibility with other measurements in the literature which use either method, and environmental trends may also reveal themselves in the relationship between these two measurements.
In Fig. 9 we show the relationship between and for our various FellWalker source extractions along with their distributions, and with a comparison to the values reported by Rigby et al. (2019) for the Inner Galaxy from the distance-limited CHIMPS sample. The much larger CHIMPS sample is illustrated as a two-dimensional hexagonal histogram to allow the point density to be seen more easily. In all cases, we find that the relationship between and is well fitted by a power law. We performed a power-law fit to the two radii types for each sample using an orthogonal distance regression555Using scipy.odr. to account for the uncertainties on both variables. The fit results are reported in Table 5. We find that the fwhires and fwlores source extractions produce essentially identical relationships, with power-law indices of and , respectively. This is unsurprising because the two extractions are very similar, differing only in the level of fragmentation that they allow (fwhires structures exist within fwlores structures).
@@
| fwhires | / pc | / pc | 0.255 | 0.007 | 0.657 | 0.010 |
| fwlores | / pc | / pc | 0.281 | 0.019 | 0.685 | 0.025 |
| fwchimps | / pc | / pc | 0.214 | 0.012 | 0.807 | 0.020 |
@@
| fwhires | / pc | / km s-1 | 0.080 | 0.016 | 0.284 | 0.027 |
| fwlores | / pc | / km s-1 | 0.040 | 0.020 | 0.405 | 0.033 |
| fwchimps | / pc | / km s-1 | 0.121 | 0.026 | 0.437 | 0.052 |
@@
| fwhires | / pc | / M⊙ | 3.558 | 0.036 | 2.010 | 0.057 |
| fwlores | / pc | / M⊙ | 3.252 | 0.055 | 1.779 | 0.081 |
| fwchimps | / pc | / M⊙ | 3.824 | 0.055 | 2.146 | 0.096 |
| Inner Galaxy | / pc | / M⊙ | 3.511 | 0.029 | 2.490 | 0.067 |
The relationship between the two radius types for the fwchimps and the Inner Galaxy samples are also non-linear and with power-law indices of and , respectively, and are consistent with each other. This suggests that there are no significant differences in the mean emission profile for sources at Galactocentric distances of 9–10 kpc compared with sources at 5–7 kpc.
The difference between the fit to the fwchimps data and the fwlores and fwhires is caused by the differences in data quality and to the FellWalker parameter setup, which we have modified in PAMS compared with CHIMPS. We note that some of the non-fitted points in Figs. 9a and b show unusually high or low values of . In cases where is much larger than , these are the archipelago sources that FellWalker identifies at low SNR. The sources with much larger than look like well-recovered sources that are diffuse and have flat emission profiles, and these are over-represented in the data due to a selection bias in FellWalker; the requirement for sources to have a minimum number of pixels above the intensity defined as the noise level prefers diffuse over compact sources. Such sources are also more likely to be the beneficiary of flux boosting effects, in which positive contributions to the emission from the noise may be represented in the data, but negative contributions will not.
Rigby et al. (2019) reported a median value of across the full CHIMPS sample, and the same value for our distance-limited Inner Galaxy sample (Fig. 9d). For fwchimps, the figure is slightly larger at 2.1. By contrast, the fwhires and fwlores median values are larger at 2.8 and 2.4, respectively, which is expected for more-sensitive data. Overall, the emission profiles of the PAMS sources are similar to those of the Inner Galaxy sample, suggesting that any differences in emission characteristics between 5–7 and 9–10 kpc are mild.
@@ 4.2.2 Scaling relationships
where is the molecular weight per hydrogen molecule, with a value of 2.8 (accounting for a 71 per cent abundance of hydrogen, 27 per cent helium, and 2 per cent metals), is the mass of a hydrogen atom, is the source distance, is the -factor calculated for 13CO (3–2) derived in Section 4.1, is the integrated intensity of 13CO (3–2) per pixel, and d is the angular area of a pixel. We adopted the Method i) values of = and cm-2 (K km s-1)-1 as our best overall estimate for resolved regions in the Inner and Outer Galaxy samples (i.e. CHIMPS and PAMS), respectively, with a factor of 1.5 uncertainty. We note that although Rigby et al. (2019) calculated clump masses for the CHIMPS sources using an LTE analysis, we apply the same derivation here for consistency with the PAMS data. We will calculate the masses of PAMS sources using LTE analysis for a comparison with the CHIMPS LTE masses and excitation conditions in a future paper.
@@
We fitted power laws to the size–linewidth and mass–radius relationships using orthogonal distance regression, as in Section 4.2.1, and present the results in Table 5. Fig. 10a shows that the size–linewidth relationships666For the uncertainty on the linewidth, we use the difference between the measured linewidth and the corrected linewidth that depends on the peak SNR of the source, which is given by . This empirical relationship was reported in Eq. 8 of Rigby et al. (2019), and accounts for the clipping of the linewidth caused by the detection thresholds used by FellWalker (and any other source-extraction algorithms that do not explicitly model emission profiles, e.g. dendrograms), though the correction factor was erroneously described as the uncertainty itself. for the fwlores and fwhires extractions show similar relationships, considering the uncertainties. Although fwlores has a slightly steeper relationship, for the most part, the fwhires sources have higher linewidths for a given size. This is a result of the greater level of fragmentation allowed in the fwhires extraction, where larger linewidths are recovered due to intra-cloud variations that are averaged out in the corresponding fwlores extraction. This is illustrated by the lines connecting the fwhires sources to their fwlores parent. For example, the NGC 7538 region catalogues contain 58 and 132 entries in the fwlores and fwhires extractions, respectively, but the bulk of the mass (95 per cent of the 13CO emission) is recovered as a single source in the fwlores catalogue, which is the parent structure of 73 of the sources (i.e. around half) featuring in the fwhires catalogue. The size–linewidth relationships for the fwchimps and CHIMPS Inner Galaxy samples are almost identical.
@@
The mass–radius relationships show more significant differences. Comparing the fwlores and fwhires source extractions, the fwhires sources have a steeper power-law index of 2.01 compared with 1.78 to fwlores, but again the difference is not significantly different when considering the uncertainties. The weak differences in the distributions correspond to a scale dependence in the density distributions of the sources; we note the greater frequency of clump-scale substructures – with sizes of pc and masses of M⊙– in the fwhires distribution. When comparing the fwchimps and Inner Galaxy samples, the power-law index is steeper in the Inner Galaxy compared with the Outer Galaxy, with a value of 2.49 for CHIMPS compared with 2.16 for PAMS, but they are consistent within the uncertainties. This consistency indicates that the mean density profile of sources in the two samples are similar.
@@
In Fig. 10, we also illustrate the relationship of (Kauffmann & Pillai, 2010) that delineates the parameter space that is empirically found to be devoid of high-mass star formation (HMSF) in nearby clouds. The original relationship was devised in terms of , with as the limit. Since we have favoured , the relationship requires modification for application to our data, and we adopt – as the average of the fitted parameters in Section 4.2.1 – and our adapted relationship is therefore . The majority of sources in the samples fitted in Fig. 10 are capable of forming high-mass stars, with 79 and 61 per cent for the fitted fwhires and fwlores samples, respectively. The proportion appears to be higher, in fact, than is the case for the CHIMPS survey, with 96 and 63 per cent of the sources from the fitted CHIMPS and fwchimps samples, respectively, satisfying the condition. We will explore why this might be the case in Section 5 for HMSF.
@@ 4.3 Galactocentric dependence
The combination of PAMS and CHIMPS allows for an expanded study of properties as a function of Galactocentric distance. The longitude coverage of CHIMPS means that the only clouds in the survey that lie outside the solar circle are at the far side of the Galaxy, with distances in the range 12–17 kpc. Consequently, those clouds are both few in number, and sample only the extreme high-column-density (and therefore high-mass) end of the underlying distributions due to Malmquist bias. The PAMS data, therefore, make an important contribution to Galactocentric trends by significantly improving the population statistics at 9–10 kpc, with much improved spatial resolution.
@@
In Fig. 11a we display the mass distributions from CHIMPS (greyscale hexagonal histogram) and PAMS (blue hexagonal histogram). For both PAMS and CHIMPS, we take the masses calculated from our factor from Section 4.1. We determined mean radial trends on a subsample of the data by first excluding all sources with peak SNRs of less than 9, as in Sections 4.2.1 and 4.2.2, and applying the mass-completeness threshold derived by Rigby et al. (2019, their equation B.1) – shown as the red dashed line – removing all sources at kpc, and all sources with M⊙. The completeness threshold should remove some of the effects of Malmquist bias. The solid orange line in Fig. 11 shows the moving average value of , with a window size of 0.1 kpc, to all sources in the subsample. It is apparent that the PAMS clumps are consistent with their Inner Galaxy counterparts, and we do not see any systematic trends over the range of probed. This is not affected by the different FellWalker parameter configurations. As far as we can tell with these data, the cloud-mass distributions in 13CO (3–2) do not vary with Galactocentric distance.
@@
In Section 4.2.2, we found slight differences in the power- law index of the mass–radius relationship between PAMS and CHIMPS clouds, indicating that the two populations may have slightly different density profiles. We therefore examine the distribution of average volume densities in Fig. 11b, calculated by:
Here we adopt the equivalent radius to ensure that the mass and radius are calculated for the same volume. We calculated average Galactic radial profiles using the same reduced sample above containing only sources satisfying the minimum mass and SNR criteria, which are, again, shown in orange. The distribution of PAMS densities for the fwhires extraction is consistent with the Inner Galaxy, whereas for fwlores the mean values are a factor of 2 lower. The latter extraction suppresses small-scale structure, and so reports cloud complexes that evidently have lower mean densities (when using for the calculation) due to the extended low-column-density envelopes that PAMS is more sensitive to. We therefore caution here that comparing distributions such as these derived from different source-extraction methods will almost certainly result in differences. When we compare like-for-like extractions, the PAMS densities within the fwchimps distribution are a factor of 4 higher than their CHIMPS counterparts, which is somewhat surprising, and we see the same offset when considering the binned radial trend. This is explainable as a distance-related bias: In the CHIMPS sample, the relationship between Galactocentric and heliocentric distances bifurcates (see Fig. 12g of Rigby et al., 2019) due to the overlapping of sources in the near and far sides of the Galactic disc in velocity. This bifurcation is visible in the distribution of mean densities in Fig. 11b at kpc, with sources at the far distance occupying the lower-density branch and vice versa for sources at the near distance. This explains this apparent trend. In the lower-right panel we demonstrate this by plotting a secondary radial trend where we limited the CHIMPS sample to sources within the same 2–4 kpc distance range that we have in our PAMS sample, and it is clear that the radial trend is consistent with that of the fwchimps extraction of PAMS. The mean densities are particularly sensitive to distance effects when using a clump-scale extraction, with more internal substructures being discernible in closer sources. Extraction techniques which prefer to recover the largest scale structures such as fwlores, or scimes (e.g Colombo et al., 2015, 2019; Duarte-Cabral et al., 2021; Rani et al., 2022) are therefore more suitable for examining Galactocentric gradients.